1. What is the percent by volume of vinegar in a sulotion that contain 3.0 ml of acetic acid dissolved in 50.5 ml water
1
Expert's answer
2021-11-09T10:28:57-0500

Solution:

solute = acetic acid (3.0 mL)

solvent = water (50.5 mL)

volume of solution = volume of solute + volume of solvent = 3.0 mL + 50.5 mL = 53.5 mL


Percent concentration by volume, (v/v)%:

(v/v)% = (volume of solute / volume of solution) × 100%

Hence,

(v/v)% = (3.0 mL / 53.5 mL) × 100% = 5.6075% = 5.6%

(v/v)% = 5.6% acetic acid


Answer: The percent by volume of vinegar is 5.6%.